OpenStudy (anonymous):

a body is falling freely under gravity.How much distance it falls during an interval of time between 1st and 2nd seconds of its motion , taking g=10?

OpenStudy (abhisar):

\(\boxed{\boxed{\boxed{\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}}}}\)

OpenStudy (anonymous):

thanks

OpenStudy (abhisar):

do u know the formula \[S=ut+\frac{ 1 }{ 2 }at ^{2}\]

OpenStudy (anonymous):

yes

OpenStudy (abhisar):

When a body falls freely, it's initial velocity is 0

OpenStudy (anonymous):

ok

OpenStudy (abhisar):

Now can u do this ?

OpenStudy (abhisar):

u have to calculate S

OpenStudy (anonymous):

yes 5*3=15?

OpenStudy (abhisar):

how u didi this ?

OpenStudy (anonymous):

1/2*10*(1.5)^2

OpenStudy (abhisar):

Between 1st and 2nd seconds means that t=2-1=1sec

OpenStudy (abhisar):

getting my points ?

OpenStudy (anonymous):

i think between 1 & 2=1.5

OpenStudy (abhisar):

NO ! interval of time between the 1st and 2nd seconds means a total duration of 2-1 = 1 sec

OpenStudy (abhisar):

Suppose second hand of a clock is at 1 sec, then it will take a time of 1 sec to reach at 2nd second....Getting it ?

OpenStudy (anonymous):

i m getting you but if i put 1.5 then my answer is right. becoz the answer is 15

OpenStudy (abhisar):

I know the answer is 15.....

OpenStudy (abhisar):

The body has travelled for a total of 2 seconds

OpenStudy (abhisar):

Now if put t=1

OpenStudy (abhisar):

It will give us the distance A

OpenStudy (abhisar):

If we put 2, it will give us the distance B

OpenStudy (abhisar):

But we have to calculate distance C

OpenStudy (abhisar):

getting it ?

OpenStudy (anonymous):

yes

OpenStudy (abhisar):

So can u calculate now ?

OpenStudy (abhisar):

10*(2)^2 - 10*(1)^2

OpenStudy (abhisar):

got it ?

OpenStudy (anonymous):

ok thanks

OpenStudy (abhisar):

\(\Huge\text{Anytime !}\) \(\huge\ddot\smile\)

OpenStudy (abhisar):

uh pretty sure, uh got this ?

OpenStudy (anonymous):

yes friend

OpenStudy (goformit100):

Hello, and A Warm Welcome to Open Study!